# Einstein from Noise: Statistical Analysis

### v2

@@ Appendix D High-dimensional signal convergence
In this section, we introduce notation and present several auxiliary results that support Theorem 4.3.
Definitions.
Recall the definition of $\tilde{\mathsf{X}}_{k}[\ell]$ from (A.21). We define:
$\sigma_{k}^{2}=\sigma^{2}\cdot\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2},$ (E.1)
@@
which corresponds to the diagonal entries of $\mathsf{\Sigma}_{k}[r,s]$ in (A.20).
Next, recall the vector $\mathsf{S}$ as defined in (A.5), and introduce the normalized vector:
$\tilde{\mathsf{S}}_{k}=\mathsf{S}/\sigma_{k},$ (E.2)
where $\sigma_{k}$ is given in (E.1). Then, by Lemma B.1, we have:
$\mathsf{\tilde{S}}_{k}|\mathsf{N}[k]\sim{\cal N}(\mathsf{\tilde{\mu}}_{k},\mathsf{\tilde{\Sigma}}_{k}),$ (E.3)
@@
$\tilde{\mu}_{k}[r]$ $\triangleq 2\sigma_{k}^{-1}\left|\mathsf{X}[k]\right|\left|\mathsf{N}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{N}_{i}}[k]-\phi_{\mathsf{X}}[k]\right)$ (E.4)
for $10\leq r\leq d-$ 1, and
$\mathsf{\tilde{\Sigma}}_{k}[r,s]$ $\triangleq\frac{\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}{\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}},$ (E.5)
@@
for $10\leq r,s\leq d-$ 1. Note in particular that normalizing $\mathsf{S}$ by $\sigma_{k}$ ensures that the diagonal entries of $\mathsf{\tilde{\Sigma}}_{k}$ are equal to one.
@@
We now present the main result of this section, which plays a key role in the proof of Theorem 4.3. To state the result, we first define the functions:
$f_{1}(r)\triangleq\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right),$ (E.6)
@@
$f_{2}(r)\triangleq\left|\mathsf{N}[k]\right|^{2}\sin^{2}\left(\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right),$ (E.7)
for $10\leq r\leq d-$ 1. Note that $f_{1}$ and $f_{2}$ correspond to the terms appearing in the expectation in the denominator and numerator of (C.8), respectively.
Lemma E.1.
Assume the template signal $x$ satisfies Assumption 4.2, and that its DFT coefficients are non-vanishing, i.e., $\mathsf{X}[k]\neq 0$, for all $10<k\leq d-$ 1. Let ${f_{1}(\hat{\mathsf{R}})}$ and ${f_{2}(\hat{\mathsf{R}})}$ be defined as in (E.6) and (E.7), with $\hat{\mathsf{R}}$ defined in (3.4). Then, as $d\to\infty$, their expected values satisfy:
$\lim_{d\to\infty}\ \frac{1}{a_{d}\cdot\left|\mathsf{X}[k]\right|}\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]=\sqrt{2}\sigma,$ (E.8)
@@
$\lim_{d\to\infty}\mathbb{E}[{f_{2}(\hat{\mathsf{R}})}]=\sigma^{2},$ (E.9)
@@
The proof of Lemma E.1 builds on Proposition D.1 and several auxiliary lemmas, which we present below. The full proof is provided in Section E.2.
E.1 Auxiliary statements
Lemma E.2.
$\lim_{d\to\infty}\sum_{\ell=0}^{d-1}\left|2\left|\mathsf{X}[\ell]\right|^{2}-\left|\mathsf{\tilde{X}}_{k}[\ell]\right|^{2}\right|=0,$ (E.10)
@@ E.1 Auxiliary statements
$\lim_{d\to\infty}\sigma_{k}^{2}=2\sigma^{2}.$ (E.11)
Proof of Lemma E.2.
From the definition of $\tilde{\mathsf{X}}_{k}[\ell]$ in (A.21), we have:
$2\left|\mathsf{X}[\ell]\right|^{2}-\left|\mathsf{\tilde{X}}_{k}[\ell]\right|^{2}=\begin{cases}2\left|\mathsf{X}[\ell]\right|^{2}&\mathsf{if}\;\ell=k,d-k,\\ \left|\mathsf{X}[\ell]\right|^{2}&\mathsf{if}\;\ell=0,d/2,\\ 0&\mathsf{otherwise}.\end{cases}$ (E.12)
According to Assumption 4.2, we have,
$\lim_{d\to\infty}\left\{{\underset{0<k\leq d-1}{\max}\left\{\left|\mathsf{X}[k]\right|\right\}}\cdot\sqrt{\log d}\right\}=0,$ (E.13)
and $\mathsf{X}[0]=0$, Recall that the template vector is normalized, i.e., $\sum_{\ell=0}^{d-1}|\mathsf{X}[\ell]|^{2}=1$, which implies $\left|\mathsf{X}[\ell]\right|^{2}\leq\left|\mathsf{X}[\ell]\right|$ for all $10\leq\ell\leq d-$ 1.
Combining (E.12)–(E.13), results,
$\lim_{d\to\infty}\sum_{\ell=0}^{d-1}\left[2\left|\mathsf{X}[\ell]\right|^{2}-\left|\mathsf{\tilde{X}}_{k}[\ell]\right|^{2}\right]=\lim_{d\to\infty}\left[\left|\mathsf{X}[0]\right|^{2}+\left|\mathsf{X}[d/2]\right|^{2}+4\left|\mathsf{X}[k]\right|^{2}\right]=0,$ (E.14)
which proves (E.10).
To establish (E.11), observe that
$\lim_{d\to\infty}\sigma_{k}^{2}=\sigma^{2}\cdot\lim_{d\to\infty}\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}=2\sigma^{2},$ (E.15)
where the final equality follows from (E.10). This completes the proof of the lemma. ∎
We now state a lemma showing that the entries of the covariance matrix $\mathsf{\tilde{\Sigma}}_{k}[r,s]$ satisfy the conditions of Proposition D.1.
Lemma E.3.
Suppose the conditions of Assumption 4.2 hold. Define,
$\rho_{\left|{r-s}\right|}\triangleq|\mathsf{\tilde{\Sigma}}_{k}\left[r,s\right]|,$ (E.16)
for $\mathsf{\tilde{\Sigma}}_{k}\left[r,s\right]$ defined in (E.5). Then, $\rho_{0}=1$, and
$\rho_{n}\log\left(n\right)\to 0.$ (E.17)
That is, the covariance matrix $\mathsf{\tilde{\Sigma}}_{k}[r,s]$ satisfies the conditions required by Proposition D.1.
Proof of Lemma E.3.
From the definition of the covariance matrix of $\mathsf{\tilde{S}}_{k}|\mathsf{N}[k]$ in (E.5), we observe that it is circulant and fully characterized by its eigenvalues $|\tilde{\mathsf{X}}_{k}[\ell]|^{2}$ (see (A.21)) for $10\leq\ell\leq d-$ 1. Due to the normalization by $\sigma_{k}$, the covariance matrix is normalized such that its diagonal entries equal one, i.e., $\rho_{0}=1$.
It remains to show that the off-diagonal elements decay sufficiently fast, namely,
$\rho_{m}\log\left(m\right)\to 0,$ (E.18)
for $m\to\infty$.
Using the definition of $\rho$ from (E.16), we can write:
$\rho_{m}=\mathsf{\tilde{\Sigma}}_{k}\left[r,r-m\right]=\frac{\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}m\right)}{\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}}.$ (E.19)
As $d\to\infty$, the denominator in (E.19) converges to 2 by Lemma E.2. The numerator corresponds to the DFT of the sequence $|\tilde{\mathsf{X}}_{k}[\ell]|^{2}$,
$\sum_{\ell=0}^{d-1}|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}m\right)=\mathcal{F}\left\{|{\tilde{\mathsf{X}}_{k}}|^{2}\right\}[m].$ (E.20)
By Lemma E.2, the following holds,
$\lim_{d,m\to\infty}\log\left(m\right)\cdot\mathcal{F}\left\{|{\tilde{\mathsf{X}}_{k}}|^{2}\right\}[m]=\lim_{d,m\to\infty}2\log\left(m\right)\cdot\mathcal{F}\left\{\left|{\mathsf{X}}\right|^{2}\right\}[m].$ (E.21)
By Assumption 4.2, the right-hand-side of (E.21) converges to zero, we have,
$\lim_{d,m\to\infty}\log\left(m\right)\cdot\mathcal{F}\left\{|{\tilde{\mathsf{X}}_{k}}|^{2}\right\}[m]=0.$ (E.22)
Combining (E.20)–(E.22) yields the desired decay condition in (E.18), thus completing the proof of the lemma. ∎
E.2 Proof of Lemma E.1
Our goal is to prove (E.8) and (E.9). By the law of total expectation, we have,
$\frac{1}{a_{d}}\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]=\frac{1}{a_{d}}\mathbb{E}\left[\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|{\mathsf{N}[k]}\right]\right],$ (E.23)
and
$\mathbb{E}[{f_{2}(\hat{\mathsf{R}})}]=\mathbb{E}\left[\mathbb{E}\left[\left.{f_{2}(\hat{\mathsf{R}})}\right|{\mathsf{N}[k]}\right]\right].$ (E.24)
Accordingly, we will prove
$\frac{\sigma_{k}}{a_{d}\cdot\left|\mathsf{X}[k]\right|}{\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]}\xrightarrow[]{\mathcal{L}^{1}}\left|\mathsf{N}[k]\right|^{2},$ (E.25)
@@
${\mathbb{E}\left[\left.{f_{2}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]}\xrightarrow[]{\mathcal{L}^{1}}\frac{1}{2}\left|\mathsf{N}[k]\right|^{2},$ (E.26)
which would yield the desired result.
To proceed, we apply Proposition D.1. Recall the definition of the vector $\tilde{\mathsf{S}}_{k}$ given in (E.2). Conditioned on $\mathsf{N}[k]$, this vector follows a Gaussian distribution with mean $\tilde{\mu}_{k}[r]$ (as defined in (E.4)) and covariance matrix $\mathsf{\tilde{\Sigma}}_{k}[r,s]$ (defined in (E.5)). We assume that both the mean and the covariance satisfy Assumption 4.2, and we assert that they also meet the criteria of Proposition D.1. Indeed, observe the following:
  1. 1.
    The mean term. By Assumption 4.2, we have $\left|\mathsf{X}[k]\right|\sqrt{\log d}\to 0$, as $d\to\infty$, for every $10\leq k\leq d-$ 1, implying that $\sqrt{\log d}\max{\left|\tilde{\mu}_{k}[r]\right|}\to 0$, where the $\left|\mathsf{N}[k]\right|$ term in $\tilde{\mu}_{k}[r]$ is finite and independent of $d$.
  2. 2.
    The covariance term. By Lemma E.3, the covariance matrix $\mathsf{\tilde{\Sigma}}_{k}[r,s]$ satisfies the conditions of Proposition D.1.
We apply Proposition D.1 and the result in (D.1) to the functions $f_{1}(\hat{\mathsf{R}})$ and $f_{2}(\hat{\mathsf{R}})$, with respect to the Gaussian vector ${\mathsf{\tilde{S}}_{k}|}\mathsf{N}[k]$ (E.3). Observe that
$\max_{0\leq r\leq d-1}{\left|\tilde{\mu}_{k}[r]\right|}=2\sigma_{k}^{-1}\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|,$ (E.27)
@@
${f_{1}(\hat{\mathsf{R}})}$ $={\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)}=\frac{\sigma_{k}}{2\left|\mathsf{X}[k]\right|}{\tilde{\mu}_{k}[\hat{\mathsf{R}}]}.$ (E.28)
Given that the assumptions of Proposition D.1 are satisfied, it follows that

### v3

@@ D.1 Notations and auxiliary results
First, we introduce notation and present several auxiliary results that support Theorem 4.3. Recall the definition of $\widetilde{\mathsf{X}}_{k}[\ell]$ from (A.21). We define:
$\sigma_{k}^{2}=\frac{\sigma^{2}}{2}\cdot\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2},$ (D.1)
@@
which corresponds to the diagonal entries of $\mathsf{\Sigma}_{k}[r,s]$ in (A.20). In addition, recall the vector $\mathsf{S}$ as defined in (A.5), and introduce the normalized vector:
$\tilde{\mathsf{S}}_{k}=\mathsf{S}/\sigma_{k},$ (D.2)
where $\sigma_{k}$ is given in (D.1). Then, by Lemma A.1, we have:
$\mathsf{\widetilde{S}}_{k}|\mathsf{N}[k]\sim{\cal N}(\mathsf{\widetilde{\mu}}_{k},\mathsf{\widetilde{\Sigma}}_{k}),$ (D.3)
@@
$\widetilde{\mu}_{k}[r]$ $\triangleq 2\sigma_{k}^{-1}\left|\mathsf{X}[k]\right|\left|\mathsf{N}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{N}_{i}}[k]-\phi_{\mathsf{X}}[k]\right)$ (D.4)
for $0\leq r\leq d-1$, and
$\mathsf{\widetilde{\Sigma}}_{k}[r,s]$ $\triangleq\frac{\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}{\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}},$ (D.5)
@@
for $0\leq r,s\leq d-1$. Note in particular that normalizing $\mathsf{S}$ by $\sigma_{k}$ ensures that the diagonal entries of $\mathsf{\widetilde{\Sigma}}_{k}$ are equal to one.
Lemma D.1.
$\lim_{d\to\infty}\sum_{\ell=0}^{d-1}\left|2\left|\mathsf{X}[\ell]\right|^{2}-\left|\mathsf{\widetilde{X}}_{k}[\ell]\right|^{2}\right|=0,$ (D.6)
and,
$\lim_{d\to\infty}\sigma_{k}^{2}=\sigma^{2}.$ (D.7)
Proof of Lemma D.1.
From the definition of $\widetilde{\mathsf{X}}_{k}[\ell]$ in (A.21), we have:
$2\left|\mathsf{X}[\ell]\right|^{2}-\left|\mathsf{\widetilde{X}}_{k}[\ell]\right|^{2}=\begin{cases}2\left|\mathsf{X}[\ell]\right|^{2}&\mathsf{if}\ell=k,d-k,\\ \left|\mathsf{X}[\ell]\right|^{2}&\mathsf{if}\ell=0,d/2,\\ 0&\mathsf{otherwise}.\end{cases}$ (D.8)
According to Assumption 4.2, we have,
$\lim_{d\to\infty}\left\{{\underset{0<k\leq d-1}{\max}\left\{\left|\mathsf{X}[k]\right|\right\}}\cdot\sqrt{\log d}\right\}=0,$ (D.9)
and $\mathsf{X}[0]=0$, Recall that the template vector is normalized, i.e., $\sum_{\ell=0}^{d-1}|\mathsf{X}[\ell]|^{2}=1$, which implies $\left|\mathsf{X}[\ell]\right|^{2}\leq\left|\mathsf{X}[\ell]\right|$ for all $0\leq\ell\leq d-1$.
Combining (D.8)–(D.9), results,
$\lim_{d\to\infty}\sum_{\ell=0}^{d-1}\left[2\left|\mathsf{X}[\ell]\right|^{2}-\left|\mathsf{\widetilde{X}}_{k}[\ell]\right|^{2}\right]=\lim_{d\to\infty}\left[\left|\mathsf{X}[0]\right|^{2}+\left|\mathsf{X}[d/2]\right|^{2}+4\left|\mathsf{X}[k]\right|^{2}\right]=0,$ (D.10)
which proves (D.6).
To establish (D.7), observe that
$\lim_{d\to\infty}\sigma_{k}^{2}=\frac{\sigma^{2}}{2}\cdot\lim_{d\to\infty}\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}=\sigma^{2},$ (D.11)
where the final equality follows from (D.6). This completes the proof of the lemma. ∎
We now state a lemma showing that the entries of the covariance matrix $\mathsf{\widetilde{\Sigma}}_{k}[r,s]$ satisfy the conditions of Proposition C.1.
Lemma D.2.
Suppose the conditions of Assumption 4.2 hold. Define,
$\rho_{\left|{r-s}\right|}\triangleq|\mathsf{\widetilde{\Sigma}}_{k}\left[r,s\right]|,$ (D.12)
for $\mathsf{\widetilde{\Sigma}}_{k}\left[r,s\right]$ defined in (D.5). Then, $\rho_{0}=1$, and
$\rho_{n}\log\left(n\right)\to 0.$ (D.13)
That is, the covariance matrix $\mathsf{\widetilde{\Sigma}}_{k}[r,s]$ satisfies the conditions required by Proposition C.1.
Proof of Lemma D.2.
From the definition of the covariance matrix of $\mathsf{\widetilde{S}}_{k}|\mathsf{N}[k]$ in (D.5), we observe that it is circulant and fully characterized by its eigenvalues $|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}$ (see (A.21)) for $0\leq\ell\leq d-1$. Due to the normalization by $\sigma_{k}$, the covariance matrix is normalized such that its diagonal entries equal one, i.e., $\rho_{0}=1$.
It remains to show that the off-diagonal elements decay sufficiently fast, namely, $\rho_{m}\log\left(m\right)\to 0$, for $m\to\infty$.
Using the definition of $\rho$ from (D.12), we can write:
$\rho_{m}=\mathsf{\widetilde{\Sigma}}_{k}\left[r,r-m\right]=\frac{\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}m\right)}{\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}}.$ (D.14)
As $d\to\infty$, the denominator in (D.14) converges to 2 by Lemma D.1. The numerator corresponds to the DFT of the sequence $|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}$,
$\sum_{\ell=0}^{d-1}|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}m\right)=\mathcal{F}\left\{|{\widetilde{\mathsf{X}}_{k}}|^{2}\right\}[m].$ (D.15)
By Lemma D.1, for each fixed $k$ and for all $m\in\{1,\ldots,d-1\}$,
$\mathcal{F}\!\big(|\widetilde{\mathsf{X}}_{k}|^{2}\big)[m]$ $=2\,\mathcal{F}\!\big(|\mathsf{X}|^{2}\big)[m]+\Delta_{k,d}[m],$ (D.16)
where the error term $\Delta_{k,d}[m]$ is supported only on the modified frequencies $\ell\in\{0,d/2,k,d-k\}$ and satisfies the uniform bound
$\sup_{1\leq m\leq d-1}\,|\Delta_{k,d}[m]|\leq|\mathsf{X}[0]|^{2}+|\mathsf{X}[d/2]|^{2}+4|\mathsf{X}[k]|^{2}\xrightarrow[d\to\infty]{}0,$ (D.17)
using Assumption 4.2(2) and $|\mathsf{X}[0]|=0$ from Assumption 4.2(3).
Moreover, recalling that $\mathsf{R}_{\mathsf{X}\mathsf{X}}={\cal F}\!\big(|\mathsf{X}|^{2}\big)$, Assumption 4.2(1) gives
$\max_{1\leq m\leq d-1}\,\log(d)\,\big|\mathcal{F}\!\big(|\mathsf{X}|^{2}\big)[m]\big|=\max_{1\leq m\leq d-1}\,\log(d)\,|\mathsf{R}_{\mathsf{X}\mathsf{X}}[m]|\xrightarrow[d\to\infty]{}0.$ (D.18)
Combining (D.16)–(D.17) with (D.18) yields
$\max_{1\leq m\leq d-1}\,\log(d)\,\big|\mathcal{F}\!\big(|\widetilde{\mathsf{X}}_{k}|^{2}\big)[m]\big|\xrightarrow[d\to\infty]{}0,$ (D.19)
and together with (D.15) this implies the desired decay condition for the off-diagonal correlations $\rho_{m}$, completing the proof. ∎
D.2 High-dimensional limits
@@
We now present a central result, which plays a key role in the proof of Theorem 4.3. To state the result, we first define the functions:
$f_{1}(r)\triangleq\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right),$ (D.20)
@@
$f_{2}(r)\triangleq\left|\mathsf{N}[k]\right|^{2}\sin^{2}\left(\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right),$ (D.21)
for $0\leq r\leq d-1$. Note that $f_{1}$ and $f_{2}$ correspond to the terms appearing in the expectation in the denominator and numerator of (B.64), respectively.
Proposition D.3 (High-dimensional limits for ${f_{1}(\hat{\mathsf{R}})}$ and ${f_{2}(\hat{\mathsf{R}})}$ ).
Assume the template signal $x$ satisfies Assumption 4.2, and that its DFT coefficients are non-vanishing, i.e., $\mathsf{X}[k]\neq 0$, for all $0<k\leq d-1$. Let ${f_{1}(\hat{\mathsf{R}})}$ and ${f_{2}(\hat{\mathsf{R}})}$ be defined as in (D.20) and (D.21), with $\hat{\mathsf{R}}$ defined in (2.3). Then, as $d\to\infty$, their expected values satisfy:
$\lim_{d\to\infty}\ \frac{1}{a_{d}\cdot\left|\mathsf{X}[k]\right|}\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]=\sigma,$ (D.22)
@@
$\lim_{d\to\infty}\mathbb{E}[{f_{2}(\hat{\mathsf{R}})}]=\frac{\sigma^{2}}{2},$ (D.23)
@@
The proof of Proposition D.3 builds on Proposition C.1 and the auxiliary Lemmas D.1D.2.
Proof of Proposition D.3.
Our goal is to prove (D.22) and (D.23). By the law of total expectation, we have,
$\frac{1}{a_{d}}\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]=\frac{1}{a_{d}}\mathbb{E}\left[\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|{\mathsf{N}[k]}\right]\right],$ (D.24)
and
$\mathbb{E}[{f_{2}(\hat{\mathsf{R}})}]=\mathbb{E}\left[\mathbb{E}\left[\left.{f_{2}(\hat{\mathsf{R}})}\right|{\mathsf{N}[k]}\right]\right].$ (D.25)
Accordingly, we will prove
$\frac{\sigma_{k}}{a_{d}\cdot\left|\mathsf{X}[k]\right|}{\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]}\xrightarrow[]{\mathcal{L}^{1}}\left|\mathsf{N}[k]\right|^{2},$ (D.26)
@@
${\mathbb{E}\left[\left.{f_{2}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]}\xrightarrow[]{\mathcal{L}^{1}}\frac{1}{2}\left|\mathsf{N}[k]\right|^{2},$ (D.27)
which would yield the desired result.
To proceed, we apply Proposition C.1. Recall the definition of the vector $\tilde{\mathsf{S}}_{k}$ given in (D.2). Conditioned on $\mathsf{N}[k]$, this vector follows a Gaussian distribution with mean $\widetilde{\mu}_{k}[r]$ (as defined in (D.4)) and covariance matrix $\mathsf{\widetilde{\Sigma}}_{k}[r,s]$ (defined in (D.5)). We assume that both the mean and the covariance satisfy Assumption 4.2, and we assert that they also meet the criteria of Proposition C.1. Indeed, observe the following:
  1. 1.
    The mean term. By Assumption 4.2, we have $\left|\mathsf{X}[k]\right|\sqrt{\log d}\to 0$, as $d\to\infty$, for every $0\leq k\leq d-1$, implying that $\sqrt{\log d}\max{\left|\widetilde{\mu}_{k}[r]\right|}\to 0$, where the $\left|\mathsf{N}[k]\right|$ term in $\widetilde{\mu}_{k}[r]$ is finite and independent of $d$.
  2. 2.
    The covariance term. By Lemma D.2, the covariance matrix $\mathsf{\widetilde{\Sigma}}_{k}[r,s]$ satisfies the conditions of Proposition C.1.
[…]
$\max_{0\leq r\leq d-1}{\left|\widetilde{\mu}_{k}[r]\right|}=2\sigma_{k}^{-1}\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|,$ (D.28)
@@
${f_{1}(\hat{\mathsf{R}})}$ $={\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)}=\frac{\sigma_{k}}{2\left|\mathsf{X}[k]\right|}{\widetilde{\mu}_{k}[\hat{\mathsf{R}}]}$. (D.29)
Given that the assumptions of Proposition C.1 are satisfied, it follows that
$\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]-\frac{\sum_{r=0}^{d-1}f_{1}(r)e^{\widetilde{\mu}_{k}[r]a_{d}}}{\sum_{r=0}^{d-1}e^{\widetilde{\mu}_{k}[r]a_{d}}}\xrightarrow[]{\mathsf{a.s.}}0,$ (D.30)
@@
$\mathbb{E}\left[\left.{f_{2}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]-\frac{\sum_{r=0}^{d-1}f_{2}(r)e^{\widetilde{\mu}_{k}[r]a_{d}}}{\sum_{r=0}^{d-1}e^{\widetilde{\mu}_{k}[r]a_{d}}}\xrightarrow[]{\mathsf{a.s.}}0.$ (D.31)
Next, we evaluate the terms at the left-hand-side of (D.30) and (D.31).
Proof of (D.26) and (D.22).
We begin by proving that
$\frac{1}{2a_{d}}\frac{\sum_{r=0}^{d-1}{\widetilde{\mu}_{k}[r]\exp\{\widetilde{\mu}_{k}[r]a_{d}}\}}{\sum_{r=0}^{d-1}{\exp\{\widetilde{\mu}_{k}[r]a_{d}}\}}\frac{\sigma_{k}^{2}}{\left|\mathsf{X}[k]\right|^{2}}\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2}.$ (D.32)
From the definition of $f_{1}(r)$, it follows that
$\sum_{r=0}^{d-1}{f_{1}(r)}=0,$ (D.33)
almost surely. Additionally, from the definition of $\widetilde{\mu}_{k}[r]$,
${\widetilde{\mu}_{k}[r]a_{d}}=2\sigma_{k}^{-1}a_{d}\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|\cos\left({\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]}\right).$ (D.34)
By Assumption 4.2, we have $a_{d}\left|\mathsf{X}[k]\right|\to 0$ as $d\to\infty$. Thus, from (D.34) and the continuous mapping theorem, we obtain
${\widetilde{\mu}_{k}[r]a_{d}}\xrightarrow[]{\mathsf{a.s.}}0.$ (D.35)
Since $\sum_{r=0}^{d-1}\widetilde{\mu}_{k}[r]=0$ almost surely (from (D.33)), and applying the continuous mapping theorem along with (D.35), we deduce that
$\frac{{\sum_{r=0}^{d-1}{\widetilde{\mu}_{k}[r]\exp\{\widetilde{\mu}_{k}[r]a_{d}}\}}}{a_{d}\sum_{r=0}^{d-1}[{\widetilde{\mu}_{k}[r]}]^{2}}\xrightarrow[]{\mathsf{a.s.}}1.$ (D.36)
Similarly, applying the continuous mapping theorem using (D.35), we get
$\frac{{\sum_{r=0}^{d-1}{\exp\{\widetilde{\mu}_{k}[r]a_{d}}\}}}{d}\xrightarrow[]{\mathsf{a.s.}}1.$ (D.37)
Next, from the definition of $\widetilde{\mu}_{k}[r]$,
$\sum_{r=0}^{d-1}[{\widetilde{\mu}_{k}[r]}]^{2}=4\sigma_{k}^{-2}\left(\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|\right)^{2}\sum_{r=0}^{d-1}\cos^{2}\left({\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]}\right).$ (D.38)
As $d\to\infty$,
$\frac{1}{d}\sum_{r=0}^{d-1}\cos^{2}\left({\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]}\right)\xrightarrow[]{\mathsf{a.s.}}\frac{1}{2}.$ (D.39)
Combining (D.38) and (D.39), we obtain
$\frac{\sigma_{k}^{2}}{2\left|\mathsf{X}[k]\right|^{2}}\sum_{r=0}^{d-1}[{\widetilde{\mu}_{k}[r]}]^{2}\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2},$ (D.40)
for $d\to\infty$. Thus, combining (D.36)–(D.40), we conclude that
$\frac{\sigma_{k}^{2}}{2\left|\mathsf{X}[k]\right|^{2}a_{d}}\frac{\sum_{r=0}^{d-1}\widetilde{\mu}_{k}[r]e^{\widetilde{\mu}_{k}[r]a_{d}}}{\sum_{r=0}^{d-1}e^{\widetilde{\mu}_{k}[r]a_{d}}}\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2}.$ (D.41)
This proves (D.32). Finally, combining (D.29), (D.30), and (D.41), we arrive at